# LimiX: Unleashing Structured-Data Modeling Capability for Generalist Intelligence

### v1

@@ 7.4 Robustness
Neural networks have been found to be vulnerable to various types of perturbations and attacks (Carlini & Wagner, 2017; Akhtar & Mian, 2018). In order to investigate the robustness of LimiX, following Hollmann et al. (2025), we conduct experiments under two types of controlled perturbations: adding uninformative features and outliers.
For uninformative features, we randomly select columns from the original dataset, shuffle each column, and concatenate them to the original dataset. For outliers, we multiply an outlier coefficient to each cell value in the original dataset with a probability of 2%. The outlier coefficient is randomly chosen between 0 and the outlier factor.
In Figure 21, we show Normalized AUC in classification tasks under the two types of perturbations. We compare LimiX with TabPFN-v2, TabICL, and CatBoost. The left figure shows that even when adding up to $90\%$ uninformative features, the normalized AUC of LimiX remains nearly unchanged. In contrast, the performance of TabICL and CatBoost drops significantly. This indicates that LimiX is much more robust than TabICL and CatBoost. The right figure shows that LimiX consistently outperforms other methods regardless of the outlier factor. In Figure 22, we show RMSE in regression tasks under perturbations. In the left figure of adding uninformative features, a similar phenomenon is observed as that in classification tasks. In the right figure of adding outliers, we find that RMSE of TabPFN-v2 rapidly increases when the outlier factor grows from 100 to 10000, while LimiX do not. This demonstrates the superior robustness of LimiX compared with TabPFN-v2 in regression tasks. Overall, LimiX exhibits superior robustness relative to the baselines.
@@ 7.5 Embedding
Pretrained models (Devlin et al., 2019; Oquab et al., 2024) are generally expected to provide effective and transferable feature representations that can be leveraged for downstream tasks. To assess the quality of embeddings extracted by LimiX, we conduct experiments on six datasets of various sample sizes and numbers of categories. We compare embeddings of LimiX with those of MLP (Goodfellow et al., 2016; Gorishniy et al., 2021), ResNet (He et al., 2016; Gorishniy et al., 2021), ModernNCA (Ye et al., 2025), TabPFN-v2 (Hollmann et al., 2025), and TabICL (Qu et al., 2025). For each model, we treat representations prior to the classification head as embeddings. Figure 23 shows t-SNE visualization of embeddings. We can see that embeddings of different categories extracted by LimiX are more separated than those extracted by other models.
To further evaluate the quality of embeddings extracted by LimiX, we additionally conduct linear probing experiments, which are widely adopted in the analysis of feature representations (Kumar et al., 2022). The experiments are conducted on BCCO-CLS. Table 22 shows that embeddings of LimiX achieve the highest average AUC and rank among the three ICL-based models. Overall, LimiX consistently outperforms baselines in both qualitative and quantitative experiments.
Figure 23: t-SNE visualization of embeddings. Different colors represent different categories. Embeddings extracted by LimiX are better separated between categories than other models.
Table 22: Results of linear probing on BCCO-CLS. LimiX outperforms the other two ICL-based models, indicating the strongest embedding capability.
@@ 7.6 Fine-tuning
Long contexts increase memory cost and can degrade optimization during fine-tuning. Therefore, it is common practice to subsample the training corpus to a manageable budget, usually via random or KNN selection (Thomas et al., 2024; Xu et al., 2025). Based on our attention-based retrieval strategy, we adopt a retrieval-guided downsampling scheme that concentrates on locally most relevant patterns within each dataset, thus improving sample efficiency and prediction performance.
@@
We compare LimiX with some representative baselines that are trained or fine-tuned on real datasets, including TabDPT (Ma et al., 2024). For the fine-tuning of TabPFN-v2, we adopt the strategy and hyperparameter configurations from its official repository [^2] https://github.com/PriorLabs/TabPFN. As shown in Tables 23 and 24, LimiX significantly outperforms the others across various metrics in most benchmarks after fine-tuning. Meanwhile, we illustrate the change of AUC for LimiX on some datasets of BCCO-CLS before and after fine-tuning. From Figure 24, a significant improvement can be observed in most cases.
@@ 7.7 Data Generation
Data generation is a challenging and meaningful task across multiple modalities since it needs to capture the joint distribution of $P(X,Y)$ compared with prediction tasks that focus on modeling $P(Y|X)$. While there is abundant literature on the generation of images (Croitoru et al., 2023), videos (Xing et al., 2024), and text (Li et al., 2024b), relatively less attention has been paid to tabular data generation. However, data generation is even more critical for tabular data due to its scarcity and possible privacy issues (Hernandez et al., 2022). As a foundation model, LimiX has the ability of tabular data generation given an unseen real dataset. A significant advantage is that it does not require additional training of generative models. Following TabPFN-v2 (Hollmann et al., 2025), firstly we conduct data generation in an iterative style. For the first column, we sample from the empirical categorical distribution if it is a categorical feature, or we use the original first column with added random noise if it is a continuous feature. For other columns, we generate $j^{th}$ column based on the j1j-1 columns of real data (treated as ${\mathbf{x}}^{ct}$) and generated data (treated as $\mathbf{x}^{\text{te}}$). We iterate this process until the last column is generated. Then for LimiX, we randomly mask a proportion of cell values of generated data and conduct missing value imputation multiple times so that we can leverage LimiX's capability of modeling the joint distribution. We conduct experiments on the following datasets: Early Stage Diabetes 2020 (Islam et al., 2019; 2020), Vertebral Column (Barreto & Neto, 2005), Seeds (Gomes Mantovani, 2015), Wine (Hardik, 2021), and Grub Damage (Vanschoren, 2014b). For evaluation, we use Trend and Shape to evaluate fidelity, which are proposed by SDMetrics [^3] https://docs.sdv.dev/sdmetrics. We also calculate the AUC of XGBoost on a hold-out test dataset using generated data or real data. From Table 25, we find that LimiX outperforms TabPFN-v2 in terms of all metrics on the five classification tasks. On the dataset of Grub Damage, the prediction performance using data generated by LimiX is even higher than using real data. The results show that LimiX is capable of capturing dependencies between features of $X$, which is brought by the pretraining strategy of mask prediction. In contrast, TabPFN-v2 is only capable of modeling $P(Y|X)$.
Table 25: Evaluation of data generation on five classification tasks. "Trend" and "Shape" are two fidelity metrics. "AUC" measures the classification performance of XGBoost using generated (or real) data. LimiX consistently outperforms TabPFN-v2 in all three metrics.
@@ 7.8 Out-of-Distribution Generalization
In real-world applications, tabular data is often subject to distribution shifts between training data and the test data encountered during deployment. For instance, a machine learning model may be trained on patient records collected from one hospital but deployed on data from a different hospital. Such distribution shifts arising from domain variation can lead to substantial degradation in model performance. This challenge is commonly referred to as the Out-of-Distribution (OOD) generalization problem, and has been the focus of extensive research in the machine learning community (Liu et al., 2021; Yu et al., 2024).
We evaluate the OOD generalization performance of various tabular models on 10 public datasets drawn from the TableShift (Gardner et al., 2023), which is a benchmark for distribution shifts in tabular data. It contains 15 binary classification tasks in total, covering a wide range of data sources, including finance, medical diagnosis, policy, etc. More details can be found in Section A.1. To ensure fair comparisons, for each dataset, if the number of training or test samples exceeds 10,000, we randomly subsample 10,000 instances. Each experiment is repeated five times and we report the average.
@@ 8 Conclusion
In this report, we introduce LimiX, an LDM that treats structured-data inputs as a joint distribution over variables and missingness, so that classification, regression, missing value imputation, data generation, and sample selection for interpretability, can all be expressed as conditional queries to a single model. Methodologically, LimiX adopts a lightweight, scalable architecture that models causal relations among variables while jointly capturing dependencies across features and samples. Meanwhile, LimiX combines masked joint-distribution pretraining with an episodic, context-conditional objective of per-dataset adaptation for the versatility in downstream tasks. The pretraining data for LimiX is generated with hierarchical SCMs via graph-aware and solvability-aware sampling. Attention-guided retrieval of LimiX further supports efficient inference-time ensemble and fine-tuning if desired. Empirically, experiments across 10 large structured-data benchmarks, spanning wide ranges of sample sizes, feature dimensions, number of classes, categorical-to-numerical feature ratios, missingness, and sample-to-feature ratios, confirm the effectiveness of LimiX. As a single model, LimiX consistently surpasses competitive baselines on various downstream tasks.
@@ A.1 Details of Datasets with Distribution Shifts
The 10 public datasets we adopt from TableShift include Voting [^4] https://electionstudies.org/, Unemployment [^5] https://www.census.gov/programs-surveys/acs, Sepsis (Reyna et al., 2020), Public Health Insurance5, Income5, Hospital Readmission (Strack et al., 2014) [^6] https://archive.ics.uci.edu/ml/datasets/Diabetes+130-US+hospitals+for+years+1999-2008, Food Stamps5, Diabetes [^7] https://www.cdc.gov/brfss/index.html, College Scorecard [^8] https://collegescorecard.ed.gov/, and ASSISTments (Metz, 2020) [^9] https://new.assistments.org. Detailed information of the benchmark TableShift we adopt in Section 7.8 can be found in Table 27.
@@ B.2.1 Sample Efficiency
Lemma B.3 (Van der Vaart (2000), Theorem 5.23; statement adapted from Qin & Risteski (2024); Li et al. (2024c)).
@@
Similar to Lemma 2 in Li et al. (2024c), we have the following lemma.
@@
Now the claim follows from Equations 7 and 9. ∎
Based on the above lemmas, we can now prove Theorem B.2. The proof follows a similar idea to that of Theorem 1 in Li et al. (2024c).

### v2

@@ 7.4 Robustness
Neural networks have been found to be vulnerable to various types of perturbations and attacks (carlini2017towards; akhtar2018threat). In order to investigate the robustness of LimiX, following hollmann2025tabpfn_nature, we conduct experiments under two types of controlled perturbations: adding uninformative features and outliers.
For uninformative features, we randomly select columns from the original dataset, shuffle each column, and concatenate them to the original dataset. For outliers, we multiply a outlier coefficient to each cell value in the original dataset with a probability of 2%. The outlier coefficient is randomly chosen between 0 and the outlier factor.
In Figure 22, we show Normalized AUC in classification tasks under the two types of perturbations. We compare LimiX with TabPFN-v2, TabICL, and CatBoost. The left figure shows that even when adding up to $90\%$ uninformative features, normalized AUC of LimiX remains nearly unchanged. In contrast, the performance of TabICL and CatBoost drops significantly. This indicates that LimiX is much more robust than TabICL and CatBoost. The right figure shows that LimiX consistently outperforms other methods regardless of the outlier factor. In Figure 23, we show RMSE in regression tasks under perturbations. In the left figure of adding uninformative features, a similar phenomenon is observed as that in classification tasks. In the right figure of adding outliers, we find that RMSE of TabPFN-v2 rapidly increases when the outlier factor grows from 100 to 10000, while LimiX do not. This demonstrates the superior robustness of LimiX-16M compared with TabPFN-v2 in regression tasks. Overall, LimiX exhibits superior robustness relative to the baselines.
@@ 7.5 Embedding
Pretrained models (devlin2019bert; oquab2023dinov2) are generally expected to provide effective and transferable feature representations that can be leveraged for downstream tasks. To assess the quality of embeddings extracted by LimiX-16M, we conduct experiments on six datasets of various sample sizes and numbers of categories. We compare embeddings of LimiX-16M with those of MLP (goodfellow2016deep; gorishniy2021revisiting), ResNet (he2016deepresnet; gorishniy2021revisiting), ModernNCA (ye2024revisitingmodernnca), TabPFN-v2 (hollmann2025tabpfn_nature), and TabICL (qu2025tabicl). For each model, we treat representations prior to the classification head as embeddings. Figure 24 shows t-SNE visualization of embeddings. We can see that embeddings of different categories extracted by LimiX-16M are more separated than those extracted by other models.
To further evaluate the quality of embeddings extracted by LimiX-16M, we additionally conduct linear probing experiments, which is widely adopted in the analysis of feature representations (kumar2022fine). The experiments are conducted on BCCO-CLS. Table 24 shows that embeddings of LimiX-16M achieve the highest average AUC and ranks among three ICL-based models. Overall, LimiX-16M consistently outperforms baselines in both qualitative and quantitative experiments.
Figure 24: t-SNE visualization of embeddings. Different colors represent different categories. Embeddings extracted by LimiX-16M are better separated between categories than other models.
Table 24: Results of linear probing on BCCO-CLS. LimiX-16M outperforms other two ICL-based models, indicating the strongest embedding capability.
@@ 7.6 Fine-tuning
Long contexts increase memory cost and can degrade optimization during fine-tuning. Therefore, it is common practice to subsample the training corpus to a manageable budget, usually via random or KNN selection (thomas2024retrieval; xu2024mixture). Based on our attention-based retrieval strategy, we adopt a retrieval-guided downsampling scheme that concentrates on locally most relevant patterns within each dataset, thus improving sample efficiency and prediction performance.
@@
We compare LimiX-16M with some representative baselines that are trained or fine-tuned on real datasets, including TabDPT (ma2024tabdpt). For the fine-tuning of TabPFN-v2, we adopt the strategy and hyperparameter configurations from its official repository [^2] https://github.com/PriorLabs/TabPFN. As shown in Tables 25 and 26, LimiX-16M significantly outperforms the others across various metrics in most benchmarks after fine-tuning. Meanwhile, we illustrate the change of AUC for LimiX-16M on some datasets of BCCO-CLS before and after fine-tuning. From Figure 25, a significant improvement can be observed in most cases.
@@ 7.7 Data Generation
Data generation is a challenging and meaningful task across multiple modalities since it needs to capture the joint distribution of $P(X,Y)$ compared with prediction tasks that focus on modeling $P(Y|X)$. While there is abundant literature on the generation of images (croitoru2023diffusion), videos (xing2024survey), and text (li2024pre), relatively less attention has been paid to tabular data generation. However, data generation is even more critical for tabular data due to its scarcity and possible privacy issues (hernandez2022synthetic). As a foundation model, LimiX models have the ability of tabular data generation given an unseen real dataset. A significant advantage is that it does not require additional training of generative models. Following TabPFN-v2 (hollmann2025tabpfn_nature), firstly we conduct data generation in an iterative style. For the first column, we sample from the empirical categorical distribution if it is a categorical feature, or we use the original first column with added random noise if it is a continuous feature. For other columns, we generate $j^{th}$ column based on the j1j-1 columns of real data (treated as ${\mathbf{x}}^{ct}$) and generated data (treated as $\mathbf{x}^{\text{te}}$). We iterate this process until the last column is generated. Then for LimiX models, we randomly mask a proportion of cell values of generated data and conduct missing value imputation for multiple times so that we can leverage LimiX's capability of modeling the joint distribution. We conduct experiments on the following datasets: Early Stage Diabetes 2020 (islam2019likelihood; early_stage_diabetes_risk_prediction_529), Vertebral Column (vertebral_column_212), Seeds (openml_seeds), Wine (openml_wine), and Grub Damage (openml_grub). For evaluation, we use Trend and Shape to evaluate fidelity, which are proposed by SDMetrics [^3] https://docs.sdv.dev/sdmetrics. We also calculate the AUC of XGBoost on a hold-out test dataset using generated data or real data. From Table 27, we find that LimiX-16M outperforms TabPFN-v2 in terms of all metrics on the five classification tasks. On the dataset of Grub Damage, the prediction performance using data generated by LimiX-16M is even higher than using real data. The results show that LimiX models are capable of capturing dependencies between features of $X$, which is brought by the pretraining strategy of mask prediction. In contrast, TabPFN-v2 is only capable of modeling $P(Y|X)$.
Table 27: Evaluation of data generation on five classification tasks. "Trend" and "Shape" are two fidelity metrics. "AUC" measures the classification performance of XGBoost using generated (or real) data. LimiX-16M consistently outperforms TabPFN-v2 in all three metrics.
@@ 7.8 Out-of-Distribution Generalization
In real-world applications, tabular data is often subject to distribution shifts between training data and the test data encountered during deployment. For instance, a machine learning model may be trained on patient records collected from one hospital but deployed on data from a different hospital. Such distribution shifts arising from domain variation can lead to substantial degradation in model performance. This challenge is commonly referred to as the Out-of-Distribution (OOD) generalization problem, and has been the focus of extensive research in the machine learning community (liu2021towards; yu2024survey).
We evaluate the OOD generalization performance of various tabular models on 10 public datasets drawn from the TableShift (gardner2023benchmarking), which is a benchmark for distribution shifts in tabular data. It contains 15 binary classification tasks in total, covering a wide range of data sources, including finance, medical diagnosis, policy, etc. More details can be found in Section A.1. To ensure fair comparisons, for each dataset, if the number of training or test samples exceeds 10,000, we randomly subsample 10,000 instances. Each experiment is repeated five times and we report the average.
@@
8 Scaling law
Recent advances in large language models (LLMs) reveal a class of empirical regularities known as scaling laws, which describe predictable relationships between model performance and key resource dimensions such as parameter count, dataset size, and compute budget (kaplan2020scaling; hoffmann2022training). These scaling laws enable researchers to forecast model behavior and guide efficient resource allocation when scaling up architectures.
However, most prior studies focus on large language models (LLMs), analyzing how loss scales with model parameters and compute under fixed or approximately fixed data distributions. Despite the growing importance of large structured-data models (LDMs), their scaling behavior remains largely unexplored.
Understanding the quantitative interaction between loss, data, and model parameters informs optimal trade-offs between data acquisition and model expansion. Motivated by this, this work systematically investigates the scaling laws for LDMs, focusing on the empirical relationships among model size, dataset size, and various performance metrics, including Log-Loss, AUC, F1, and accuracy for classification, as well as RMSE-Loss and $R^{2}$ for regression tasks. Our primary focus is on evaluating these metrics on downstream tasks, as they directly reflect the model's practical effectiveness and its ability to generalize to real-world applications. By analyzing performance on tasks like classification and regression, we aim to identify the optimal balance between model size, data, and task-specific performance.
For the experimental setup, we construct a series of models with parameter counts ranging from 1.05M to 16.53M (and up to 1.6G in extrapolated fits), trained on datasets spanning $1\times 10^{11}$ to $1.7\times 10^{12}$ tokens. Each configuration maintains consistent architecture and optimization settings, isolating the effects of data and parameter scaling. The downstream benchmark we used is BCCO-CLS and BCCO-REG. For this experiment, the evaluation was performed without the retrieval procedure, which may lead to minor deviations in performance compared to the results reported in the experimental section.
8.1 Scaling with Dataset Size and Model Parameters
Following previous findings in large language models (LLMs), we assume that both the loss and performance metrics of our model follow a power-law relationship with respect to computational resources, including dataset size and model parameter count. This assumption reflects the empirical regularity that as data volume or model capacity increases, the model’s loss tends to decrease while its predictive performance improves in a predictable manner. Formally, we hypothesize that each evaluation metric $M$ (e.g., Log-Loss, RMSE-Loss, AUC, ACC, or $R^{2}$) can be expressed as a power-law function of the resource variable $N$:
$M=a\cdot N^{c}+b,$ (5)
where $a$, $b$, and $c$ are constants to be determined. To estimate these parameters, we employ an ordinary least squares (OLS) regression over measurements collected across multiple scales of $N$. For each metric and task configuration, the fitted parameters and the coefficient of determination ($R^{2}$) are reported to assess the strength and consistency of the scaling behavior. This procedure enables systematic quantification of how both loss reduction and performance improvement scale with resource growth, facilitating direct comparisons between different model sizes and data regimes.
Figure 26 illustrates the empirical loss scaling with dataset size across different model configurations. All models achieve highly consistent fits with $R^{2}>0.99$, confirming the robustness of the scaling law. As dataset size increases, the loss decreases rapidly at smaller scales and then transitions to a more gradual improvement trend. Larger models achieve consistently lower loss values and exhibit smoother, more stable scaling behavior as additional data are introduced.
(a) Log-Loss versus dataset size on BCCO-CLS
(b) RMSE-Loss versus dataset size on BCCO-REG
Figure 26: Loss versus dataset size for models of different parameter counts. Dotted lines indicate fitted power-law relations.
Furthermore, we extend our analysis to examine how downstream task loss and performance scale with model parameters when unconstrained by dataset size or computational budget. As shown in Figure 27, multiple performance metrics exhibit clear dependencies on the model parameter count $N$. For the classification task, the empirical fits achieve $R^{2}>0.9$ for Log-Loss, AUC, and ACC, confirming the robustness of the scaling law, while the fit for F1 yields a slightly lower $R^{2}$ of 0.68. For the regression task, both RMSE-Loss and the $R^{2}$ score exhibit strong correlations with model size, with empirical fits achieving $R^{2}>0.9$.
The empirical fits are reported below, with only those achieving $R^{2}>0.9$ shown:
[…]
Loss decays sharply with increasing $N$, while AUC, ACC and F1 improve gradually, indicating that while error convergence is achieved relatively early, fine-grained discriminative improvements require substantially larger models.
(a) Log-Loss and performance versus parameter count on BCCO-CLS
(b) RMSE-Loss and performance versus parameter count on BCCO-REG
Figure 27: Joint scaling of Loss, AUC, F1, ACC, and $R^{2}$ with respect to parameter count. Real and fitted trends are shown as solid and dashed lines, respectively.
8.2 Conclusions and Insights
This study provides the first systematic characterization of scaling laws in LDMs. Our key findings are summarized as follows:
Loss versus Dataset Size: both the classification (BCCO-CLS) and regression (BCCO-REG) tasks exhibit a consistent power-law relationship between loss and dataset size. When the token count increases from approximately $10^{11}$ to $10^{13}$, the loss decreases systematically across all model configurations. Smaller models (e.g., 1.05M parameters) display a sharper initial decline in loss, while larger models (e.g., 16.53M parameters) achieve overall lower loss levels and maintain more stable improvement trends as data volume grows. The fitting results demonstrate exceptionally high explanatory power ($R^{2}>0.99$), confirming the robustness of the scaling behavior. These observations suggest that increasing the dataset size consistently enhances model performance, with improvement rates varying according to model capacity.
Loss and Performance versus Model Parameters As the number of model parameters increases from 1.05M to approximately one billion, the loss decreases markedly while overall predictive performance improves. In classification tasks, loss declines rapidly at smaller scales and stabilizes at larger ones, while AUC and ACC values continue to increase consistently. Similarly, in regression tasks, RMSE decreases and the coefficient of determination ($R^{2}$) rises with model size, both following strong power-law trends ($R^{2}>0.98$). These results suggest that scaling up the model enhances predictive stability and generalization. However, the diminishing rate of improvement in loss relative to performance metrics implies that additional parameters primarily contribute to refined discriminative capability rather than further error convergence.
Overall, the scaling behavior of the LDM follows a predictable power-law trend analogous to that observed in large language models. Both increased dataset size and model capacity yield systematic and highly correlated reductions in loss ($R^{2}>0.99$). Nevertheless, the returns on scaling diminish progressively, as larger models require substantially greater data and computational resources to achieve marginal gains in performance. These results highlight the importance of balancing dataset expansion and model scaling to optimize the trade-off between predictive accuracy and computational efficiency.
@@ 9 Conclusion
In this work, we present LimiX, a unified family of our LDM series, and release its first two variants, LimiX-16M and LimiX-2M. LimiX models treat structured-data inputs as a joint distribution over variables and missingness, so that classification, regression, missing value imputation, data generation, and sample selection for interpretability, can all be expressed as conditional queries to a single model. Methodologically, LimiX models adopt a lightweight, scalable architecture that models causal relations among variables while jointly capturing dependencies across features and samples. Meanwhile, LimiX models combine masked joint-distribution pretraining with an episodic, context-conditional objective of per-dataset adaptation for the versatility in downstream tasks. The pretraining data for LimiX models is generated with hierarchical SCMs via graph-aware and solvability-aware sampling. Attention-guided retrieval of LimiX models further supports efficient inference-time ensemble and fine-tuning if desired. Empirically, experiments across 11 large structured-data benchmarks, spanning wide ranges of sample sizes, feature dimensions, number of classes, categorical-to-numerical feature ratios, missingness, and sample-to-feature ratios, confirm the effectiveness of LimiX models . As a single model, LimiX-16M consistently surpasses competitive baselines on various downstream tasks. We present the first systematic investigation of scaling laws in LDMs, revealing a predictable power-law relationship between scale and performance, analogous to that observed in large language models.
@@ A.1 Details of Datasets with Distribution Shifts
The 10 public datasets we adopt from TableShift include Voting [^4] https://electionstudies.org/, Unemployment [^5] https://www.census.gov/programs-surveys/acs, Sepsis (reyna2020early), Public Health Insurance5, Income5, Hospital Readmission (strack2014impact) [^6] https://archive.ics.uci.edu/ml/datasets/Diabetes+130-US+hospitals+for+years+1999-2008, Food Stamps5, Diabetes [^7] https://www.cdc.gov/brfss/index.html, College Scorecard [^8] https://collegescorecard.ed.gov/, and ASSISTments (metz2020assistments) [^9] https://new.assistments.org. Detailed information of the benchmark TableShift we adopt in Section 7.8 can be found in Table 29.
@@ B.2.1 Sample Efficiency
Lemma B.3 (van2000asymptotic, Theorem 5.23; statement adapted from qin2024fit; li2024promises).
@@
Similar to Lemma 2 in li2024promises, we have the following lemma.
@@
Now the claim follows from Equations 9 and 11. ∎
Based on the above lemmas, we can now prove Theorem B.2. The proof follows a similar idea to that of Theorem 1 in li2024promises.